A building contractor can purchase a truck from her local Toyota dealer for GH¢100,000. The Toyota
dealer can also lease the truck for GH¢24,400 per year over five years. The truck has an expected life of
five years. The truck is expected to be worth GH¢10,000 in five years and the contractor intends to buy it
at fair market value at that time. If the contractor wants to purchase the truck, she must borrow the money
from IBK Bank at a current rate of 20%. The contractor’s tax rate is 30% for simplicity, assume straightline depreciation and the entire lease rentals are tax deductible. The contractor will incur GH¢5,000 in
maintaining and insuring the truck. Would leasing be a better option?
Answer #1

In the present case the comparision is required between Yearly Lease Rental for Truch Vs Purchase of Truck through Bank Finance with the interest rate of 20%. The Cash Flows are as Under.

Purchase Price of Truck is GH¢100,000

Year Lease Rentals Bank Interest Post Tax Lease Rentals Post Tax Bank Interest
1 24,400

100,000*20% =

20,000

24,400* (1-30%) =

17,080

20,000* (1-30%) =

14,000

2 24,400

100,000*20% =

20,000

24,400* (1-30%) =

17,080

20,000* (1-30%) =

14,000

3 24,400

100,000*20% =

20,000

24,400* (1-30%) =

17,080

20,000* (1-30%) =

14,000

4 24,400

100,000*20% =

20,000

24,400* (1-30%) =

17,080

20,000* (1-30%) =

14,000

5 24,400

100,000*20% =

20,000

24,400* (1-30%) =

17,080

20,000* (1-30%) =

14,000

TOTAL 85,400 70,000

In this case Tax rate is 30% hence Post Tax Lease Rentals amounts for 5 years is GH¢ 85,400 where as Interest payments to Bank is GH¢ 70,000.

Futher GH¢5,000 inmaintaining and insuring the truck will remain same in both the option of Lease and Purchase. And in case of Purchase the Terminal Value of the truck at the end of 5 years is GH¢ 10,000.

Hence taking all into considerations Purchase option with Bank loan at 20% is better.